OpenStudy (anonymous):

At the picnic, there is a plate of 3 dill pickles, 2 bread and butter pickles, and 1 sweet pickle. What is the probability that Shawn will randomly pick a dill pickle and then a sweet pickle to eat?

OpenStudy (anonymous):

\[\frac{1}{2}\times \frac{1}{5}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

clear or no?

OpenStudy (anonymous):

yea its 1/10 thank you again